Professor Geoffrey Fox

Northeast Parallel Architectures Center 111 College Place, 3-217 CST Syracuse University

Phone: (315) 443-2163 FAX: (315) 443-1973 Email: gcf@npac.syr.edu

February 15, 1996

Employee Benefits Skytop Syracuse University

To whom it may concern,

This is to confirm that Paul Coddington will be based in New York City on a long-term basis while working for the Northeast Parallel Architectures Center (NPAC).

Dr Coddington will be working part of his time on a federally-funded contract to develop a high-performance software library, and the rest of his time will involve working with companies in the New York City and Hudson Valley areas, as part of NPAC’s InfoMall technology transfer program.

Yours sincerely,

Geoffrey Fox Director of NPAC Professor of Physics and Computer Science